OpenStudy (anonymous):

(y-2/y) diveded by (y+2/y) what would you do?

OpenStudy (anonymous):

multiply top and bottom by 7 to clear the compound fraction

OpenStudy (anonymous):

I would divide them. Try it yourself and post what you get

OpenStudy (anonymous):

(y^2-2)/(y^2+2)

OpenStudy (amistre64):

id start by crying, followed by some sort of chocolate icecream

OpenStudy (anonymous):

\[\frac{y-\frac{2}{y}}{y+\frac{2}{y}}\] \[\frac{y-\frac{2}{y}}{y+\frac{2}{y}}\times \frac{y}{y}\] \[\frac{y^2-2}{x^2+2}\]

OpenStudy (anonymous):

what color is that? pepto bismol?

OpenStudy (aravindg):

hi guys plzzz help me after answering this question

OpenStudy (amistre64):

\[\frac{y-2/y}{y+2/y}=N\] \[y-2/y=N(y+2/y)\] \[y^2-2=N(y^2+2)\] \[\frac{y^2-2}{y^2+2}=N\]